# Naeural AI OS - Decentralized ubiquitous computing MLOps execution engine

### v4

@@ Abstract
Over the past few years, ubiquitous, or pervasive computing has gained popularity as the primary approach for a wide range of applications, including enterprise-grade systems, consumer applications, and gaming systems. Ubiquitous computing refers to the integration of computing technologies into everyday objects and environments, creating a network of interconnected devices that can communicate with each other and with humans. By using ubiquitous computing technologies, communities can become more connected and efficient, with members able to communicate and collaborate more easily. This enabled interconnectedness and collaboration can lead to a more successful and sustainable community. The spread of ubiquitous computing, however, has emphasized the importance of automated learning and smart applications in general. Even though there have been significant strides in Artificial Intelligence and Deep Learning, large scale adoption has been hesitant due to mounting pressure on expensive and highly complex cloud numerical-compute infrastructures. Adopting, and even developing, practical machine learning systems can come with prohibitive costs, not only in terms of complex infrastructures but also of solid expertise in Data Science and Machine Learning. In this paper we present an innovative approach for low-code development and deployment of end-to-end AI cooperative application pipelines. We address infrastructure allocation, costs, and secure job distribution in a fully decentralized global cooperative community based on tokenized economics.
@@ 1 Introduction
1.2 The Naeural Edge Protocol vision
@@ 3.3 Secure execution with Plugin API
It is important to note that all these approaches can be accessed and utilized through external endpoints. This can be done either via a JavaScript/TypeScript-based SDK or through a Python client that utilizes our Naeural E2 PyE2 package.
@@ 3.5 Distributed decentralized execution
To enhance the adoption of the ecosystem by the public, software development kits (SDKs) have been developed. These SDKs support multiple programming languages and frameworks, such as Python and NodeJS, significantly expanding integration options and improving the user experience. By using the SDKs, users gain access to the Naeural Network through a partially functional paradigm, which offers callbacks for manipulating payloads received from Naeural Nodes and provides simple yet powerful features for communication with the network. This empowers users to connect to the network, establish new end-to-end pipelines or join existing ones, create predefined or custom jobs within the pipeline, and receive notifications and heartbeats, facilitating seamless interaction with the ecosystem.
The Naeural SDKs are intended to be used as a cohesive integration tool between the AI processing network and any existing or new software applications. They are designed to automate and manage the lifecycle of requests made on the Naeural Network. Naeural SDKs come in several flavors targeting either the Python or Node.js developers’ communities.
@@
3.5.1 The Naeural SDKs
The Naeural Node.js SDK consists of two main components. The first component aids in project setup and maintenance, handling tasks such as code generation and message validation using YAML schemas. The second component is responsible for runtime operations, serving as an intermediary for communication with the network and an example of it can be seen in example 2
To support developers in integrating network functionality into their solutions, a Command Line Interface (CLI) tool is provided. This tool assists in constructing the necessary schemas for communicating with deployed pipelines in the Naeural Network. During project development or maintenance, developers can use the CLI to ensure that the command definitions for the pipelines are synchronized with the network.
The application runtime incorporates a package tailored for a specific programming language, which is meant to be seamlessly integrated into the business process services of the consuming application. The implementation details may vary depending on the programming language, but generally involve callbacks which trigger internal events. The responses obtained from these events are then analyzed and converted into data structures that are specific to the programming language in use. This allows developers to design and create business objects that align with the requirements and objectives of the application.
To incorporate Naeural into a project, there are two approaches available. The first option is to execute a series of CLI commands using the Naeural Node.js SDK. This will retrieve the necessary AI pipeline message schemas specifically designed for Node.js applications. Alternatively, for Python development teams, the Naeural PyE2 SDK package can be installed, eliminating the need for CLI commands. Once this prerequisite is fulfilled, the imported libraries within the project will handle all validation and transformations between the application’s business objects and the expected message formats of the Naeural Nodes.
@@ 3.7 Naeural Rapid Application Development
The Naeural Ecosystem envisions a future where AI-powered application development is accessible to individuals, regardless of their familiarity with programming languages. To achieve this, we are creating tools that enable both developers and non-developers to effortlessly articulate their business requirements and translate them into instructions compatible with the Naeural SDKs.
@@ 4.2.2 Stock prediction
The communication between the spreadsheet and the API can be achieved through a RESTful architecture or in real-time using a duplex channel such as websockets. The API utilizes the Naeural Node.js SDK and taps into the predictive models published within the network to generate the results that are then relayed back to the user of the spreadsheet. For a more detailed depiction of the underlying processes, refer to the swimlane diagram provided in Figure4.
Figure 4: Naeural network integration in a Google Sheets plugin
@@ 4.2.4 Use-case: Anomaly Detection
4.3 Video Anonymization - Motionmask™
Motionmask™  a partner app developed using the open source stack of Naeural Edge Protocol, illustrates the integration of the Naeural Network within the backend infrastructure of a traditional Web2 service. The application serves as a web-based tool for anonymizing videos and ensuring compliance with data protection regulations through the provision of an user-friendly interface for identifying and masking specific individuals or vehicles. In order to have a clear view, the swimlane diagram from Figure5 describes the underlying processes in more detail.
Figure 5: Naeural Network integration in Motionmask™information flows
At the core of Motionmask™ lie very powerful AI detection models. These models require a lot of computing resources which, since Motionmask™ is an online available tool, would have to be provided solely by Motionmask™ itself. This would not be possible and efficient without the Naeural Network because it would require for Motionmask™ to have an infrastructure small enough to avoid wastage, but also big enough to reliably support surges in user requests.
The interface, which operates within a web browser, communicates with the backend by making RESTful API calls. The backend, in turn, manages all the application business logic such as handling user creation and deletion, registering requests for video anonymization and handling file uploads. The integration of the Naeural Network is done within a dedicated internal service that connects to the provided MQTT broker and communicates with the nearest Naeural Node.
Progress is tracked by witnessing the events emitted by the Naeural Node which describe the state of the running pipeline. Upon completion of processing, the backend services publish an event that is propagated back to the user interface, and the resultant output is presented to the user.
The payment for services provided by the Motionmask™ application can be completed using both traditional payment processing services and cryptocurrency. Specifically, payments can be made through traditional payment processing channels with FIAT currency or through the transmission of tokens to the Motionmask™ smart contract.

### v6

@@ Abstract
Over the past few years, ubiquitous, or pervasive computing has gained popularity as the primary approach for a wide range of applications, including enterprise-grade systems, consumer applications, and gaming systems. Ubiquitous computing refers to the integration of computing technologies into everyday objects and environments, creating a network of interconnected devices that can communicate with each other and with humans. By using ubiquitous computing technologies, communities can become more connected and efficient, with members able to communicate and collaborate more easily. This enabled interconnectedness and collaboration can lead to a more successful and sustainable community. The spread of ubiquitous computing, however, has emphasized the importance of automated learning and smart applications in general. Even though there have been significant strides in Artificial Intelligence and Deep Learning, large scale adoption has been hesitant due to mounting pressure on expensive and highly complex cloud numerical-compute infrastructures. Adopting, and even developing, practical machine learning systems can come with prohibitive costs, not only in terms of complex infrastructures but also of solid expertise in Data Science and Machine Learning. In this paper we present an innovative approach [^1] This project incorporates open-source components developed with the support of financing grant SMIS 156084 ”ReDeN - Decentralized Network for Neural Processing”, provided by the Romanian Competitiveness Operational Programme. We extend our gratitude for this support, which has been instrumental in advancing our work and enabling us to share these resources with the international community. for low-code development and deployment of end-to-end AI cooperative application pipelines. We address infrastructure allocation, costs, and secure job distribution in a fully decentralized global cooperative community based on tokenized economics.
@@ 1 Introduction
1.2 The Ratio1 Protocol vision
@@
However, democratizing AI is not just a matter of technological efficiency; it is a fundamental shift towards accessibility, affordability, and inclusion in AI-driven innovation. Today, AI advancements are often concentrated in the hands of large corporations with the resources to develop, deploy, and commercialize cutting-edge models. This centralized model creates barriers to entry, limiting opportunities for individuals, small businesses, and underserved communities to participate in the AI economy.
Our approach aims to eliminate these barriers by creating an open, trustless ecosystem where AI processing power is distributed across a decentralized network. This brings several key societal benefits:
  • Accessibility: By leveraging distributed computing, AI services can reach users who lack access to high-performance infrastructure. Whether it’s a student in a rural area needing access to AI-powered learning tools or a healthcare provider in a developing region requiring AI diagnostics, our vision ensures that advanced AI applications are no longer a privilege of the few.
  • Affordability: Decentralized AI deployment reduces reliance on expensive, proprietary cloud solutions. By optimizing resource sharing and allowing anyone with computing power to contribute, we enable cost-effective AI services that are affordable for startups, independent developers, and organizations with limited budgets.
  • Inclusion in Underserved Markets: Many industries and regions suffer from a lack of AI integration due to prohibitive costs or infrastructure constraints. Our model fosters AI innovation in emerging markets by lowering the financial and technological barriers to adoption. This can lead to breakthroughs in sectors such as agriculture, education, and healthcare, where AI has the potential to drive significant positive change.
@@ 3.3 Secure execution with Plugin API
It is important to note that all these approaches can be accessed and utilized through external endpoints. This can be done either via a Python client that utilizes our Naeural E2 PyE2 package or other similar approaches regardless of the programming language.
@@ 3.5 Distributed decentralized execution
To enhance the adoption of the ecosystem by the public, software development kits (SDKs) have been developed. These SDKs support multiple programming languages and frameworks, such as Python, significantly expanding integration options and improving the user experience. By using the SDKs, users gain access to the Naeural Network through a partially functional paradigm, which offers callbacks for manipulating payloads received from Naeural Nodes and provides simple yet powerful features for communication with the network. This empowers users to connect to the network, establish new end-to-end pipelines or join existing ones, create predefined or custom jobs within the pipeline, and receive notifications and heartbeats, facilitating seamless interaction with the ecosystem.
The Naeural SDKs are intended to be used as a cohesive integration tool between the AI processing network and any existing or new software applications. They are designed to automate and manage the lifecycle of requests made on the Naeural Network. Naeural SDKs come in several flavors targeting either the Python or other developers’ communities.
@@
3.5.1 The Naeural SDK
To support developers in integrating network functionality into their solutions, a Command Line Interface (CLI) tool is provided - namely nepctl. This tool assists in various tasks such as obtaining available online nodes within Naeural Network or configure the SDK installation.
Note that each individual SDK will have a private key generated upon the first installation on the target computer. This private key allows the SDK client to connect to the Naeural network and even pair the local SDK with certain nodes by explicitly adding the SDK public key in the edge node allowed list.
The application runtime incorporates a package tailored for a specific programming language, which is meant to be seamlessly integrated into the business process services of the consuming application. The implementation details may vary depending on the programming language - currently supporting Python - but generally involve callbacks which trigger internal events. The responses obtained from these events are then analyzed and converted into data structures that are specific to the programming language in use. This allows developers to design and create business objects that align with the requirements and objectives of the application.
For Python development teams, the Naeural PyE2 SDK package can be installed and once this prerequisite is fulfilled, the imported libraries within the project will handle all validation and transformations between the application’s business objects and the expected message formats of the Naeural Nodes.
@@ 3.7 Naeural Rapid Application Development
The Naeural Ecosystem envisions a future where AI-powered application development is accessible to individuals, regardless of their familiarity with programming languages. To achieve this, we are creating tools that enable both developers and non-developers to effortlessly articulate their business requirements and translate them into instructions compatible with the Naeural SDK.
@@ 4.2.2 Stock prediction
The web-API that serves the predictions in this example relies on the Naeural Node internal FastAPI-based plugins that support both RESTful architecture as well as statefulness.